OpenStudy (anonymous):

A chemist added an excess of sodium sulfate to a solution of a soluble barium compound to precipitate all of the barium ion barium sulfate, BaSO4. How many grams of barium ion are in a 403 mg sample of the precipitate? What is the mass percentage of barium in the compound? So far, I have the balanced equation: Na2+SO4 + Ba = BaSO4 + 2Na

OpenStudy (chmvijay):

first find molar mass of BaSO4 let us assume its X X mass of BaSO4 contains =6.023*10^23 atom/ions of Ba 403 mg of BaSO4 contains = 6.023*10^23 *403 mg / X =